OpenStudy (anonymous):

STATS: if two people are playing independent games, separately (like solitaire) but taking turns, and person A has a .4 prob of winning and person B has a .7 prob of winning, what is the prob that person B will win first?

OpenStudy (anonymous):

First, the probability that B wins on their first turn is: P(A loses first turn)*P(B wins first turn) = 0.6 * 0.7 = 0.42 right?

OpenStudy (anonymous):

Basically, you have to calculate an infinite series... Familiar with those?

OpenStudy (anonymous):

It works out to about 51.2% chance that B wins, if I'm not mistaken...

OpenStudy (anonymous):

If you want me to explain this, tell me the part that is confusing you.